OpenStudy (anonymous):

Put x-y=-4 into slope intercept form? WILL MEDAL AND FAN ;)

OpenStudy (alyssa_xo):

solve for y recall slope-intercept form \(y=mx+b\)

OpenStudy (anonymous):

is the answer y = -4 + x?

OpenStudy (alyssa_xo):

I'd write it as y=x-4

OpenStudy (anonymous):

But its the same thing, right

OpenStudy (alyssa_xo):

they are, but conventionally the x comes first. that's just how things are

OpenStudy (alyssa_xo):

well mx, but in this case m=1

OpenStudy (anonymous):

Thanks @alyssa_xo

OpenStudy (alyssa_xo):

anytime bbygurl

OpenStudy (anonymous):

its actually y=x+4, you were close, but you missed distributing the negative sign

OpenStudy (alyssa_xo):

whoops

OpenStudy (anonymous):

So i graphed it wrong baaaaaaaaah

OpenStudy (anonymous):

@jeweaver6895 change y = 2 - 1/2x to slope intercept form?

OpenStudy (anonymous):

that already is in slope intercept form, the only thing you could change would be to move around the variables a little so you get y=-1/2x+2